What is the pH of a buffer with 0.15M HN3 and 0.20 M LiN3? Ka= 4.5x10-4
Answer #1

[Salt] = [ Li N3] = 0.20M [Acid) = [H N3] = 0.15 M ka = 4.5 x 104 pka = -log ka =-log (4-5X16+) pH = pka + log [salt) pH = -l